Convert to Farenheit
Use the formula:
F = 9/5 (fraction).C + 32
96 degrees C = _____ degrees F

Answer by checkley77(12844)

F = 9/5 (fraction).C + 32
F=9/5(96)+32
F=9*96/5+32
F=854/5+32
F=172.8+32
F=204.8 ANS.